OpenStudy (anonymous):

m^3/8(m^-1/4 +9m) simplify

OpenStudy (anonymous):

m^3/8(1/m(4+9m) =m^3 / (8/4m+9m^2) =m^3 * (4m+9m^2 /8) =4m^4 + 9m^5 / 8